# Multi-task learning for molecular electronic structure approaching coupled-cluster accuracy

### v1

@@ Abstract
Machine learning (ML) plays an important role in quantum chemistry, providing fast-to-evaluate predictive models for various properties of molecules. However, as most existing ML models for molecular electronic properties use density function theory (DFT) databases as the ground truth in training, their prediction accuracy cannot go beyond the DFT. In this work, we developed a unified ML method for electronic structures of organic molecules using the gold-standard CCSD(T) calculations as training data. Tested on hydrocarbon molecules, our model outperforms the DFT with the widely-used B3LYP functional in both computation costs and prediction accuracy of various quantum chemical properties. We apply the model to aromatic compounds and semiconducting polymers on both ground state and excited state properties, demonstrating its accuracy and generalization capability to complex systems that are hard to calculate using CCSD(T)-level methods.
@@ I Introduction
Computational methods for molecular systems play an essential role in physics, chemistry, and materials science, providing predictive tools for revealing underlying mechanisms of physical phenomena and accelerating materials design [1, 2]. Among various types of computational methods, quantum chemistry calculations of electronic structure are usually the bottleneck, limiting the computational speed and scalability [3]. In recent years, machine learning (ML) methods have been successfully applied to accelerate molecular dynamics simulations and to improve their accuracy in many application scenarios [4, 5, 6]. Particularly, ML inter-atomic potential can predict energy and force of molecular systems with significantly lower computational costs compared to quantum chemistry methods [5, 6, 7, 8]. Indeed, recent advances in universal ML potential enable large-scale molecular dynamics simulation with the complexity of realistic physical systems [9, 10, 11, 12, 13].
Besides energy and force, electronic properties that explicitly involve the electron degrees of freedom are also essential in molecular simulation [14]. In the past few years, ML methods have also been extended to electronic structure of molecules, providing predictions for various electronic properties such as electric multipole moments [15, 16, 17], electron population [18], excited states properties [19, 20], as well as the electronic band structure of condensed matter [21, 22]. Most of these methods take the density functional theory (DFT) results as the training data, using neural networks (NN) to fit the single-configurational representation (either the Kohn-Sham Hamiltonian or molecular orbitals) of the DFT calculations [15, 21, 19, 23]. Along with the rapid advances of machine learning techniques, the NN predictions match the DFT results increasingly well, approaching the chemical accuracy (1 kcal/mol) [16, 9]. However, as a mean-field level theory, DFT calculations themselves induce a systematic error, which is usually several times larger than the chemical accuracy [24], limiting the overall accuracy of the ML model trained on DFT datasets.
In comparison, the correlated wavefunction method CCSD(T) is considered the gold-standard in quantum chemistry [25]. It provides high accuracy predictions on various molecular properties. Unfortunately, the computational cost of CCSD(T) calculations has a rather unfavorable scaling relationship with system size. Hence, it can only handle small molecules with up to hundreds of electrons. This urges the combination of CCSD(T) and ML methods, which can potentially have both high accuracy and low computational cost. However, previous methods that directly fit the single-configurational representation of the DFT calculations cannot be applied with the CCSD(T) training data. This is because CCSD(T) does not provide either a Kohn-Sham (KS) Hamiltonian or single-body molecular orbitals due to the many-body quantum entanglement nature of its representation.
@@
In this work, we develop a unified multi-task machine learning method for molecular electronic structures that takes the CCSD(T) calculations as training data. Our method incorporates the equivariant graph neural network (EGNN) [5, 6, 7, 26], where vectors and tensors are involved in the message-passing step. Using hydrocarbon organic molecules as a testbed, our method predicts molecular energy within chemical accuracy as compared with both CCSD(T) calculation and experiments, and predicts electric dipole and quadrupole moments, atomic charge, bond order, energy gap, and electric polarizability with better accuracy than one of the most widely used DFT functional, namely B3LYP [27]. Our trained model shows robust generalization from small molecules in the training dataset (molecular weight $<100$) to larger molecules and even semiconducting polymers (molecular weight up to several thousands). Systematically predicting multiple electronic properties using a single model with local DFT computational speed, our method provides a high-performance tool for computational organic chemistry and a promising framework for ML electronic structure calculations.
@@ II Results
Here we briefly describe the computational workflow and model architecture of our EGNN method. Given an input atomic configuration, our methods output an effective single-body Hamiltonian matrix that predicts involved quantum chemical properties, as shown in Fig. 1a. The workflow consists of the input layer, convolutional layer, and output layer.
The input layer takes atomic configurations as input, including the information of atomic numbers $(Z_{1},Z_{2},\cdots,Z_{n})$ and atomic coordinates $(\vec{r}_{1},\vec{r}_{2},\cdots,\vec{r}_{n})$ of a $n$-atom system. A molecular graph is constructed, where atoms are mapped to graph nodes, while bonds between atoms (neighboring atoms within a cut-off radius $r_{\rm cut}=2$) are mapped to graph edges. The atomic configuration is then encoded into the node features $\mathbf{x}_{i,\rm in}$ and edge features $\mathbf{f}_{ij,\rm in}$ (see Methods A for details). The electron wavefunction is described in the medium-sized cc-pVDZ basis set $\{|\phi_{i,\alpha}\rangle\}$ [28], where $i$ is the index of atom and $\alpha$ is the index of basis function. We use the ORCA quantum chemistry program package [29] (the Orbital Integrator block in Fig. 1a) to calculate the overlap matrix $S_{i\alpha,j\beta}\equiv\langle\phi_{i,\alpha}|\phi_{j,\beta}\rangle$ and a single-body effective Hamiltonian $F_{i\alpha,j\beta}$ (where $i\alpha$ is the row index and $j\beta$ is the colunm index). The $F_{i\alpha,j\beta}$ matrix is obtained by a DFT calculation using a fast-to-evaluate local functional BP86 [30]. The Lowdin-symmetrized KS Hamiltonian [31] is then obtained as
@@
where the last term is an identity shift to account for the many-body energy term (see Methods A for details). Note that $F^{\prime}$ is a local DFT-level effective Hamiltonian, meaning that its accuracy is relatively low. We will use $F^{\prime}$ as the starting point of our ML model, as explained below.
@@
The convolutional layer uses EGNN implemented by the e3nn package [26] to transform the input node features $\mathbf{x}_{i,\rm in}$ and edge features $\mathbf{f}_{ij,\rm in}$ to the output node features $\mathbf{x}_{i,\rm out}$ and edge features $\mathbf{f}_{ij,\rm out}$. The EGNN includes a series of linear transformation (Self-Interaction block), activation function (Non-Linearity block), and graph convolution (Tensor Product and Concatenation blocks) layers, as shown in Fig. 1b. Details on the numerical form and dimension of each block are elaborated in Methods B. The output $\mathbf{f}_{ij,\rm in}$ and $\mathbf{x}_{i,\rm out}$ encode equivariant features of atoms and bonds as well as their atomic environment.
@@
Then, we construct an equivariant ML correction Hamiltonian $V^{\theta}$ from the output features $\mathbf{x}_{i,\rm out}$ and $\mathbf{f}_{ij,\rm out}$ as follow:
@@
$V_{i\alpha,j\beta}^{\theta}=\begin{cases}V_{\rm node}(\mathbf{x}_{i,\rm out})_{\alpha,\beta}&\text{if $i=j$}\\ \frac{V_{\rm edge}(\mathbf{f}_{ij,\rm out})_{\alpha,\beta}+V_{\rm edge}(\mathbf{f}_{ji,\rm out})_{\beta,\alpha}}{2}&\text{if $i\neq j$}\end{cases},$ (2)
@@
where $V_{\rm node}(\mathbf{x}_{i,\rm out})$ rearranges node features to a $N_{i}\times N_{i}$ symmetric matrix, and $V_{\rm edge}(\mathbf{f}_{ij,\rm out})$ rearranges edge features to a $N_{i}\times N_{j}$ matrix ($N_{i},N_{j}$ are the numbers of basis functions of the atom $i,j$). Note that the output matrices $V^{\theta}$ are Hermitian and equivariant under rotation according to the transformation rule of the basis set $\{|\phi_{i,\alpha}\rangle\}$ (see Methods B for details).
Finally, the total effective Hamiltonian of the system ${\bf H}^{\rm eff}={\bf F}^{\prime}+{\bf V}^{\theta}$ is obtained by adding the correction term $V^{\theta}$ to the BP86 Kohn-Sham Hamiltonian ${\bf F}^{\prime}$ to account for the non-local exchange-correlation effects. An effective electronic structure of the molecule is obtained by solving the eigenvalue equations of $H_{i\alpha,j\beta}^{\rm eff}$:
$\sum_{j,\beta}H_{i\alpha,j\beta}^{\rm eff}c_{j,\beta}^{n}=\epsilon_{n}c_{i,\alpha}^{n},$ (3)
@@
where $\epsilon_{n}$ gives the $n$th energy levels, and $c^{n}_{i,\alpha}$ gives the corresponding molecular orbitals through basis expansion $|\psi_{n}\rangle=\sum_{i,\alpha}\tilde{c}^{n}_{i,\alpha}|\phi_{i,\alpha}\rangle$ with $\tilde{c}^{n}=S^{-1/2}c^{n}$.
We then use the obtained energy levels and molecular orbitals to evaluate a series of ground state properties $O_{g}$ according to the rules of quantum mechanics:
$O_{g}^{\rm EGNN}=f_{O_{g}}(\{\epsilon_{n}\},\{\mathbf{c}^{n}\}),\quad O_{g}=E,\vec{p},\mathbf{Q},C_{i},B_{ij},$ (4)
@@
where properties $O_{g}$ goes through the ground state energy ($E$), electric dipole ($\vec{p}$) and quadrupole ($\mathbf{Q}$) moments, Mulliken atomic charge [32] of each atom $C_{i}$, and Mayer bond order [33] of each pair of atoms $B_{ij}$. We also evaluate the excited state properties $O_{e}$ including the energy gap (first excitation energy, $E_{\rm g}$) and static electric polarizability $\alpha$:
$O_{e}^{\rm EGNN}=f_{O_{e}}(\{\epsilon_{n}\},\{\mathbf{c}^{n}\},\mathbf{G},\mathbf{T}),\qquad O_{e}=E_{\rm g},\mathbf{\alpha}.$ (5)
@@
In principle, the ground state electronic structure does not contain the information of excited states. Therefore, we use the EGNN-output correction terms $\mathbf{G}$ (energy gap correction) and $\mathbf{T}$ (dielectric screening matrix) to account for the excited states information. The function forms of $f_{O_{g}}$ and $f_{O_{e}}$ are elaborated in Methods C.
@@
Figure 2: Benchmark of the model performance on the testing dataset. (a) Mean-square error loss of different quantities as a function of training dataset size. (b) Computational costs of different methods plot against number of electrons. The computational costs is measured as the calculation time on a Intel Xeon Platinum 8260 CPU node on the MIT SuperCloud [34] with sufficient memory for all calculations. (c) Prediction errors of the energy (E per atom, reference to separate atoms), electric dipole moment (p), electric quadrupole moment (Q), Mulliken atomic charge (C), Mayer bond order (B), energy gap (1st excitation energy, $E_{\rm g}$), and static electric polarizability ($\alpha$, a.u. means atomic unit). The blue and orange bars represent the root-mean-square error of EGNN and B3LYP calculations on configurations with each chemical formula. We plot a representative atomic configuration of each chemical formula below the horizontal axis.
@@
Table 1: Benchmark of EGNN model’s error in predicting different quantum chemical properties on the in-domain (ID) testing dataset and out-of-domain (OOD) validation dataset compared to the B3LYP density functional theory calculations. The B3LYP calculation uses def2-SVP basis set implemented by The ORCA quantum chemistry program package [29].
Properties Unit EGNN (ID) B3LYP (ID) EGNN (OOD) B3LYP (OOD)
Energy (per atom) kcal/mol $\mathbf{0.11}$ 2.19 $\mathbf{0.10}$ 2.41
Electric Dipole Debye $\mathbf{0.012}$ 0.024 $\mathbf{0.015}$ 0.023
@@
Mulliken Atomic Charge e $\mathbf{0.038}$ 0.189 $\mathbf{0.033}$ 0.195
Mayer Bond Order $\mathbf{0.023}$ 0.053 $\mathbf{0.017}$ 0.034
1st-Excitation Energy eV $\mathbf{0.26}$ 0.58 $\mathbf{0.31}$ 0.63
Polarizability a.u. $\mathbf{1.85}$ 2.22 $\mathbf{3.91}$ 4.32
Figure 3: Experimental validation of the EGNN’s predictions on gas phase standard enthalpy of formation of aromatic hydrocarbon molecules. The EGNN predictions (blue bars, right axis) and experimental values from Ref. [35] (breen bars, right axis) are compared for 11 molecules (see SI Table II for details). The difference between the EGNN method and experimental values are shown by the orange line (left axis), and the experimental uncertainty is shown by the red line (left axis).
The multi-task learning is implemented by minimizing the loss function $L_{\rm Total}$ constructed as follows:
$L_{\rm Total}$ $=l_{V}+\sum_{O\in\{O_{g}\},\{O_{e}\}}l_{O},$ (6)
$l_{V}$ $=\frac{w_{V}}{N_{\rm basis}^{2}}\sum_{i\alpha,j\beta}(V^{\theta}_{i\alpha,j\beta})^{2},$
@@
Here for each property $O$, $l_{O}$ is the the mean-square error (MSE) loss between $O^{\rm EGNN}$ and $O^{\rm label}$, the EGNN predictions (Eq. (4)(5)) and coupled-cluster labels in the training dataset, respectively. Meanwhile, $l_{V}$ is a regularization that penalizes large correction matrix $V_{\theta}$, and $N_{\rm basis}$ is the total number of basis functions in the molecule. The weights $w_{V}$ and $w_{O}$ are hyperparameters whose values are listed in supplementary information (SI) section I. The weights are chosen according to the relative importance of each property (e.g., the largest weight is assigned to $w_{E}$ due to the importance of energy). Minimizing $L_{\rm Total}$ requires the back-propagation through Eq. (3) (i.e., calculating $\partial\epsilon_{n}/\partial\mathbf{H}^{\rm eff}$ and $\partial\mathbf{c}^{n}/\partial\mathbf{H}^{\rm eff}$), which is numerically unstable [36]. To overcome this issue, we derive customized back-propagation schemes for each property using perturbation theory in quantum mechanics (see SI section S1 for details).
@@
Atomic configurations of molecules in our dataset are generated by the workflow shown in Fig. 1c. First, 85 hydrocarbon molecule structures are collected from the PubChem database [37]. Molecular dynamics (MD) simulation with TeaNet interatomic potential [5, 6] is then performed for each molecule structure to sample an ensemble of atomic configurations. Then, the CCSD(T) calculations are used to calculate the ground-state properties of sampled configurations, and the EOM-CCSD [38] calculations are used to calculate the excited-state properties. The EGNN model is trained on small-molecules training dataset (training domain, Fig. 1c) and tested on both small molecules in the training dataset (in-domain validation) and larger molecules outside the training dataset (out-of-domain validation). Details about training and testing dataset and training hyperparameters are elaborated in SI section S2. The model’s generalization capability from small molecules to large molecules is essential for its usefulness on complex systems where coupled cluster calculations cannot be implemented on the current computational platforms, due to their formidable computational costs.

### v2

@@ Abstract
Machine learning (ML) plays an important role in quantum chemistry, providing fast-to-evaluate predictive models for various properties of molecules. However, most existing ML models for molecular electronic properties use density functional theory (DFT) databases as ground truth in training, and their prediction accuracy cannot surpass that of DFT. In this work, we developed a unified ML method for electronic structures of organic molecules using the gold-standard CCSD(T) calculations as training data. Tested on hydrocarbon molecules, our model outperforms DFT with the widely-used hybrid and double hybrid functionals in computational costs and prediction accuracy of various quantum chemical properties. As case studies, we apply the model to aromatic compounds and semiconducting polymers on both ground state and excited state properties, demonstrating its accuracy and generalization capability to complex systems that are hard to calculate using CCSD(T)-level methods.
@@ I Introduction
Computational methods for molecular and condensed matter systems play essential roles in physics, chemistry, and materials science, which can reveal underlying mechanisms of diverse physical phenomena and accelerate materials design [1, 2]. Among various types of computational methods, quantum chemistry calculations of electronic structure are usually the bottleneck, limiting the computational speed and scalability [3]. In recent years, machine learning (ML) methods have been successfully applied to accelerate molecular dynamics simulations and to improve their accuracy in many application scenarios [4, 5, 6]. Particularly, ML inter-atomic potential can predict energy and force of molecular systems with significantly lower computational costs compared to quantum chemistry methods [5, 6, 7, 8]. Indeed, recent advances in universal ML potential enable large-scale molecular dynamics simulation with the complexity of realistic physical systems [9, 10, 11, 12, 13]. Besides ML inter-atomic potential, rapid advances also appear in another promising direction, namely the ML density functional, which focuses on further improving the energy prediction towards the chemical accuracy (1 kcal/mol) [14, 15, 16, 17].
Besides energy and force, other electronic properties that explicitly involve the electron degrees of freedom are also essential in molecular simulations [18]. In the past few years, ML methods have also been extended to electronic structure of molecules, predicting various electronic properties such as electric multipole moments [19, 20, 21], electron population [22], excited states properties [23, 24], as well as the electronic band structure of condensed matter [25, 26]. Most of these methods take the density functional theory (DFT) results as the training data, using neural networks (NN) to fit the single-configurational representation (either the Kohn-Sham Hamiltonian or molecular orbitals) of the DFT calculations [19, 25, 23, 27]. Along with the rapid advances of ML techniques, the NN predictions match the DFT results increasingly well, approaching the chemical accuracy [20, 9]. However, as a mean-field level theory, DFT calculations themselves induce a systematic error, which is usually several times larger than the chemical accuracy [28], limiting the overall accuracy of the ML model trained on DFT datasets.
@@
In comparison, the correlated wavefunction method CCSD(T) is considered the gold-standard in quantum chemistry [29]. It provides high accuracy predictions on various molecular properties. Unfortunately, the computational cost of CCSD(T) calculations has a rather unfavorable scaling relationship with system size. Hence, it can only handle small molecules with up to hundreds of electrons. This urges the combination of CCSD(T) and ML methods, which can potentially have both high accuracy and low computational cost. However, above-mentioned ML methods that directly fit the single-configurational representation of the DFT calculations cannot be applied with the CCSD(T) training data. This is because CCSD(T) does not provide either a Kohn-Sham (KS) Hamiltonian or single-body molecular orbitals due to the many-body quantum entanglement nature of its representation.
In this work, we develop a unified multi-task ML method for molecular electronic structures. Instead of focusing solely on energy, our method also provides accurate predictions for various electronic properties; compared with ML models trained on DFT datasets, our method learns from CCSD(T)-accuracy training data. The method incorporates the equivariant graph neural network (EGNN) [5, 6, 7, 30], where vectors and tensors are involved in the message-passing step. Using hydrocarbon organic molecules as a testbed, our method predicts molecular energy within chemical accuracy as compared with both CCSD(T) calculation and experiments, and predicts various properties, including electric dipole and quadrupole moments, atomic charge, bond order, energy gap, and electric polarizability with better accuracy than B3LYP, one of the most widely used hybrid DFT functional [31]. Our trained model shows robust generalization capability from small molecules in the training dataset (molecular weight $<100$) to larger molecules and even semiconducting polymers (molecular weight up to several thousands). Systematically predicting multiple electronic properties using a single model with local DFT computational speed, the method provides a high-performance tool for computational chemistry and a promising framework for ML electronic structure calculations.
@@ II Results
II.1 Theory and Model Architecture
In this section, we briefly describe the theoretical background and model architecture of our EGNN method. Basically, we use a NN to simulate the non-local exchange-correlation interactions of a many-body system. Then, a physics-informed approach is used to predict multiple properties from the output of a single NN.
Computational Workflow. Given an input atomic configuration, our methods output an effective single-body Hamiltonian matrix that predicts quantum chemical properties, as shown in Fig. 1a. The workflow consists of the input layer, the convolutional layer, and the output layer.
The input layer takes atomic configurations as input, including the information of atomic numbers $(Z_{1},Z_{2},\cdots,Z_{n})$ and atomic coordinates $(\vec{r}_{1},\vec{r}_{2},\cdots,\vec{r}_{n})$ of a $n$-atom system. A molecular graph is constructed, where atoms are mapped to graph nodes, while bonds between atoms (neighboring atoms within a cut-off radius $r_{\rm cut}=2$) are mapped to graph edges. The atomic configuration is then encoded into the node features $\mathbf{x}_{I,\rm in}$ for atom information and edge features $\mathbf{f}_{IJ,\rm in}$ for bond information (see Methods A for details). The electron wavefunction is represented using an atomic orbital basis set $\{|\phi_{I,\mu}\rangle\}$ [32], where $I$ is the index of atom and $\mu$ is the index of atomic orbital basis function. Then, a quantum chemistry calculation [33] (the Orbital Integrator block in Fig. 1a) is used to evaluate single-body effective Hamiltonian $F_{I\mu,J\nu}$ and the overlap matrix $S_{I\mu,J\nu}\equiv\langle\phi_{I,\mu}|\phi_{J,\nu}\rangle$ in the non-orthogonal atomic-orbital representation (where $I\mu$ is the row index and $J\nu$ is the colunm index). The $F_{I\mu,J\nu}$ matrix is obtained by a fast-to-evaluate single-configurational method such as local DFT [34]. The Lowdin-symmetrized KS Hamiltonian [35] is then obtained as
@@ II.1 Theory and Model Architecture
where the last term is an identity shift to account for the many-body energy term (see Methods A for details). Note that $\bf{F}^{\prime}$ is a local DFT-level effective Hamiltonian, meaning that it is easy and fast to compute, but its accuracy is relatively low. We will use $\bf{F}^{\prime}$ as the starting point of our ML model, and the total effective Hamiltonian of the system ${\bf H}^{\rm eff}={\bf F}^{\prime}+{\bf V}^{\theta}$ is obtained by adding the ML correction term ${\bf V}^{\theta}$. As ${\bf F}^{\prime}$ is obtained from a local DFT calculation, it only contains local exchange-correlation contribution, and the correction term ${\bf V}^{\theta}$ would account for the non-local exchange-correlation effects. Generally, the non-local exchange-correlation effects can be incorporated in CCSD(T) calculations. However, as mentioned before, the computational costs of CCSD(T) methods are formidably high for large system. The essence of our ML method is to obtain the non-local exchange-correlation effects from a NN, whose computational cost scales only linearly with system size.
@@
To obtain the ML correction term, we build a neural network model to predict ${\bf V}^{\theta}$. The EGNN framework is employed for the convolutional layer because of its outstanding performance in predicting molecular properties [7]. The convolutional layer transforms the input node features $\mathbf{x}_{I,\rm in}$ and edge features $\mathbf{f}_{IJ,\rm in}$ to the output node features $\mathbf{x}_{I,\rm out}$ and edge features $\mathbf{f}_{IJ,\rm out}$. The EGNN includes a series of linear transformation (Self-Interaction block), activation function (Non-Linearity block), and graph convolution (Tensor Product and Concatenation blocks) layers, as shown in Fig. 1b. Details on the numerical form and dimension of each block are elaborated in Methods B. The output $\mathbf{f}_{IJ,\rm out}$ and $\mathbf{x}_{I,\rm out}$ encode equivariant features of atoms and bonds as well as their atomic environment.
@@
Then, we construct an equivariant ML correction Hamiltonian $\bf{V}^{\theta}$ from the output features $\mathbf{x}_{I,\rm out}$ and $\mathbf{f}_{IJ,\rm out}$ as follow:
@@
$V_{I\mu,J\nu}^{\theta}=\begin{cases}\left[V_{\rm node}(\mathbf{x}_{I,\rm out})\right]_{\mu,\nu}&\text{if $I=J$}\\ \frac{1}{2}\left[V_{\rm edge}(\mathbf{f}_{IJ,\rm out})\right]_{\mu,\nu}+\frac{1}{2}\left[V_{\rm edge}(\mathbf{f}_{JI,\rm out})\right]_{\nu,\mu}&\text{if $I\neq J$}\end{cases}$ (2)
@@
where $V_{\rm node}(\mathbf{x}_{I,\rm out})$ is a $N_{I}\times N_{I}$ symmetric matrix rearranged from node features $\mathbf{x}_{I,\rm out}$, while $V_{\rm edge}(\mathbf{f}_{IJ,\rm out})$ is a $N_{I}\times N_{J}$ matrix obtained from edge features $\mathbf{f}_{JI,\rm out}$. Here $N_{I},N_{J}$ are the numbers of basis functions of the atom $I,J$. Note that the output matrices $\bf{V}^{\theta}$ are Hermitian and equivariant under rotation according to the transformation rule of the basis set $\{|\phi_{I,\mu}\rangle\}$ (see Methods B for details).
An effective electronic structure of the molecule is obtained by solving the eigenvalue equations of the total Hamiltonian $H_{I\mu,J\nu}^{\rm eff}$:
$\sum_{J,\nu}H_{I\mu,J\nu}^{\rm eff}c_{J,\nu}^{i}=\epsilon_{i}c_{I,\mu}^{i},$ (3)
@@
where $\epsilon_{i}$ gives the $i$-th energy levels, and $c^{i}_{I,\mu}$ gives the corresponding molecular orbitals through basis expansion $|\psi_{i}\rangle=\sum_{I,\mu}\tilde{c}^{i}_{I,\mu}|\phi_{I,\mu}\rangle$ with $\tilde{c}^{i}=S^{-1/2}c^{i}$.
Multiple Learning Tasks. Our scheme aims to predict multiple observable molecular properties (more than just energy). In order to achieve reduced computational costs, we do not include information about the entire electronic Hilbert space as learning targets. The energy levels and molecular orbitals are used to evaluate a series of ground state properties $O_{g}$ according to the rules of quantum mechanics:
$O_{g}^{\rm EGNN}=f_{O_{g}}(\{\epsilon_{i}\},\{\mathbf{c}^{i}\}),\quad O_{g}=E,\vec{p},\mathbf{Q},C_{I},B_{IJ},$ (4)
@@
where properties $O_{g}$ goes through the ground state energy ($E$), electric dipole ($\vec{p}$) and quadrupole ($\mathbf{Q}$) moments, Mulliken atomic charge [36] of each atom $C_{I}$, and Mayer bond order [37] of each pair of atoms $B_{IJ}$. We also evaluate the energy gap (first excitation energy, $E_{\rm g}$) and static electric polarizability $\alpha$:
$E_{g}^{\rm EGNN}$ $=f_{E_{g}}(\{\epsilon_{i}\},\{\mathbf{c}^{i}\},\mathbf{G}),$ (5)
$\alpha^{\rm EGNN}$ $=f_{\alpha}(\{\epsilon_{i}\},\{\mathbf{c}^{i}\},\mathbf{T}).$
@@
In principle, the ground state electronic structure does not contain the information of the energy gap and electric polarizability. Therefore, we use the EGNN-output correction terms $\mathbf{G}$ (energy gap correction) and $\mathbf{T}$ (dielectric screening matrix) to account for the excited states information and the linear response information, respectively. The function forms of $f_{O_{g}}$, $f_{E_{g}}$, and $f_{\alpha}$ are elaborated in Methods C. Note that these properties are all derived from the underlying electronic structure, so they are internally related. Therefore, multi-task learning methods can utilize these relations to enhance the model’s generalization capability.
@@
Figure 2: Benchmark of the model performance on the testing dataset. (a) Testing root-mean-square errors (RMSE) of different quantities as a function of training dataset size. (b) Computational costs of different methods plot against number of electrons. The computational costs is measured as the calculation time (node hour) on a single Intel Xeon Platinum 8260 CPU node with 48 cores on the MIT SuperCloud [38] with sufficient memory for all calculations. The scaling deviates from the theoretical asymptotic scaling, e.g., $N^{7}$ for CCSD(T), because the parallelization efficiency is higher for larger molecules. In principle, the $N^{7}$ scaling for CCSD(T) would appear in the large $N$ limit. (c) Prediction RMSE of the energy ($E$ per atom, reference to separate atoms), electric dipole moment ($\vec{p}$), electric quadrupole moment ($\bf{Q}$), Mulliken atomic charge ($C$), Mayer bond order ($B$), energy gap (1st excitation energy, $E_{\rm g}$), and static electric polarizability ($\alpha$, a.u. means atomic unit). Our EGNN method is compared with the B3LYP hybrid functional, DSD-PBEP86 double hybrid functional [39], DM21 ML functional [14], and AIQM1 ML potential [13, 40]. A representative atomic configuration of each chemical formula is plotted for illustration.
@@
Table 1: Benchmark of EGNN model’s RMSE in predicting different quantum chemical properties on the in-domain (ID) testing dataset and out-of-domain (OOD) validation dataset (written as ID RMSE/OOD RMSE in the table).
RMSE (ID/OOD) Unit Hybrid Double hybrid DM21 AIQM1 EGNN (ours)
Energy (per atom) kcal/mol 2.20/2.41 0.94/1.20 0.22/0.11 0.13/0.06 0.11/0.10
Electric Dipole 10-2 Debye 1.27/1.20 0.71/0.70 0.78/0.88 0.63/0.81
@@
Mulliken Atomic Charge e 0.19/0.20 0.04/0.05 0.05/0.04 0.04/0.03
Mayer Bond Order 0.05/0.03 0.04/0.02 0.06/0.03 0.02/0.02
1st-Excitation Energy eV 0.59/0.63 3.71/3.26 1.71/1.47 0.26/0.31
Polarizability a.u. 2.22/4.32 4.74/8.05 1.85/3.91
The multi-task learning is implemented by minimizing the total loss function $L_{\rm Total}$ constructed as follows:
$L_{\rm Total}$ $=\sum_{O\in\{O_{g},E_{g},\alpha\}}l_{O}+l_{V},$ (6)
@@
Here for each property $O$, $l_{O}$ is the the mean-square error (MSE) loss between $O^{\rm EGNN}$ and $O^{\rm label}$, the EGNN predictions (Eq. (4)(5)) and coupled-cluster labels in the training dataset, respectively. Meanwhile, $l_{V}$ is a regularization that penalizes large correction matrix $\bf{V}_{\theta}$, and $N_{\rm basis}$ is the total number of basis functions in the molecule. The weights $w_{V}$ and $w_{O}$ are hyperparameters whose values are listed in supplementary information (SI) section I. The weights are chosen to balance the training tasks so that the training errors of all tasks decrease to satisfactory levels. Minimizing $L_{\rm Total}$ requires the back-propagation through Eq. (3) (i.e., calculating $\partial\epsilon_{i}/\partial\mathbf{H}^{\rm eff}$ and $\partial\mathbf{c}^{i}/\partial\mathbf{H}^{\rm eff}$), which is numerically unstable [41]. To overcome this issue, we derive customized back-propagation schemes for each property using perturbation theory in quantum mechanics (see SI section S1 for details), giving
$\nabla_{\theta}\epsilon_{i}$ $=(\mathbf{c}^{i})^{\dagger}(\nabla_{\theta}\bf{V}^{\theta})\mathbf{c}^{i}$ (7)
$\nabla_{\theta}\mathbf{c}^{i}$ $=\sum_{p\neq i}\frac{(\mathbf{c}^{p})^{\dagger}(\nabla_{\theta}\mathbf{V}^{\theta})\mathbf{c}^{i}}{\epsilon_{i}-\epsilon_{p}}\mathbf{c}^{p}.$
When evaluating the gradients of properties in Eqs. (4) and (5) using the chain rule, terms that analytically cancel each other are removed in numerical evaluation, making the scheme numerically stable.
@@
Software Realization. Atomic configurations of molecules in our training dataset are generated by the workflow shown in Fig. 1c. First, 85 hydrocarbon molecule structures are collected from the PubChem database [42]. Molecular dynamics (MD) simulation with TeaNet interatomic potential [5, 6] is then performed for each molecule structure to sample an ensemble of atomic configurations. Then, the CCSD(T) calculations are used to calculate the ground-state properties of sampled configurations, and the EOM-CCSD [43] calculations are used to calculate the excited-state properties. The $\bf{S}$ and $\bf{F}$ matrices are evaluated by the ORCA quantum chemistry program package [33] with the fast-to-evaluate BP86 local density functional [34] and the medium-sized cc-pVDZ basis set [32]. The EGNN model based on the e3nn package [30] is trained on small-molecules training dataset (training domain, Fig. 1c) and tested on both small molecules in the training dataset (in-domain validation) and larger molecules outside the training dataset (out-of-domain validation). Details about training and testing dataset and training hyperparameters are elaborated in SI section S2.
II.2 Model Performance and Applications
In this section, we benchmark the performance of our EGNN model and display potential applications of the model in systems of practical importance. We will make direct comparisons with experimental results wherever possible, and the results suggest the outstanding generalization capability and predicting power of our approach.